Calculating EVA Brewster Company manufactures elderberry wine. Last year, Brewster earned operating income of $194,000 after...

  1. Calculating EVA

    Brewster Company manufactures elderberry wine. Last year, Brewster earned operating income of $194,000 after income taxes. Capital employed equaled $2.6 million. Brewster is 40 percent equity and 60 percent 10-year bonds paying 7 percent interest. Brewster’s marginal tax rate is 40 percent. The company is considered a fairly risky investment and probably commands a 13-point premium above the 5 percent rate on long-term Treasury bonds.

    Jonathan Brewster’s aunts, Abby and Martha, have just retired, and Brewster is the new CEO of Brewster Company. He would like to improve EVA for the company. Compute EVA under each of the following independent scenarios that Brewster is considering.

    Required:

    Use a spreadsheet to perform your calculations and round all interim and percentage figures to four decimal places. If the EVA is negative, enter your answer as a negative amount.

    1. No changes are made; calculate EVA using the original data.

    $

    2. Sugar will be used to replace another natural ingredient (atomic number 33) in the elderberry wine. This should not affect costs but will begin to affect the market assessment of Brewster Company, bringing the premium above long-term Treasury bills to 11 percent the first year and 8 percent the second year. Calculate revised EVA for both years.

    EVA
    Year 1 $
    Year 2 $

    3. Brewster is considering expanding but needs additional capital. The company could borrow money, but it is considering selling more common stock, which would increase equity to 80 percent of total financing. Total capital employed would be $4,000,000. The new after-tax operating income would be $380,000. Using the original data, calculate EVA. Then, recalculate EVA assuming the materials substitution described in Requirement 2. New after-tax income will be $380,000, and in Year 1, the premium will be 11 percent above the long-term Treasury rate. In Year 2, it will be 8 percent above the long-term Treasury rate. (Hint: You will calculate three EVAs for this requirement.)

    EVA
    Year 1 $
    Year 1 (11% premium) $
    Year 2 (8% premium) $
Answer #1

Part 1)

We will have to calculate the weighted average cost of capital to determine the EVA as follows:

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 60%*7%*(1-40%) + 40%*(5% + 13%) = 9.72%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 194,000 - 2,600,000*9.72% = -$58,720

Part 2)

The EVA for both the years is calculated as below:

Year 1:

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 60%*7%*(1-40%) + 40%*(5% + 11%) = 8.92%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 194,000 - 2,600,000*8.92% = -37,920

_____

Year 2:

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 60%*7%*(1-40%) + 40%*(5% + 8%) = 7.72%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 194,000 - 2,600,000*7.72% = -$6,720

_____

Tabular Representation:

EVA

Year 1

-$37,920

Year 2

-$6,720

_______

Part 3)

The original EVA for Year 1 with revised values is calculated as below:

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 30%*7%*(1-40%) + 70%*(5% + 13%) = 13.86%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 380,000 - 4,000,000*13.86% = -$174,400

_____

Year 1 (Requirement 2):

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 30%*7%*(1-40%) + 70%*(5% + 11%) = 12%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 380,000 - 4,000,000*12% = -$100,000

_____

Year 2 (Requirement 2):

WACC = Weight of Debt*After-Tax Cost of Debt + Weight of Equity*Cost of Equity = 30%*7%*(1-40%) + 70%*(5% + 8%) = 10.4%

Now, we can calculate EVA as below:

EVA = Net Operating Income - Capital Employed*WACC = 380,000 - 4,000,000*10.4% =- $36,000

_____

Tabular Representation:

EVA

Year 1

-$174,400

Year 1 (11% Premium)

-$100,000

Year 2 (8% Premium)

-$36,000